Amanda will receive a $20 000 annual payment from a family trust fund. This will continue until she reached age 40. She is currently 25 years old and will receive one payment per a year. Inflation is 5% and the interest rate of the trust fund is 6% compounded monthly. What is the value of the trust fund today?  
Answer #1

Number of years she receives the 20,000 annual payment =15 years

effective annual rate with monthly compounding = (1+0.06/12)^12 -1 = 0.0617 = 6.17%

Value of the trust fund today =PV(rate,nper,pmt) in excel =PV(0.0617,15,20000) = 192,133.14

Value of the trust fund today = $192,133.14